# thanks: Kavli IPMU Fellow. Corresponding author

### v1

@@ Abstract
The study of ultralight dark matter helps constrain the lower bound on minimally coupled dark matter models. The granular structure of ultralight dark matter density fields produces metric perturbations which have been identified as a potentially interesting probe of this model. For dark matter masses $m\gtrsim 10^{-17}\,\mathrm{eV}$ these perturbations would fluctuate on timescales similar to observational timescales. In this paper, we estimate the expected time delay these fluctuations would generate in simulated pulsar signals. We simulate arrays of mock pulsars in a fluctuating granular density field. We calculate the expected Shapiro time delay and gravitational redshift and compare analytical estimates with the results of simulations. Finally, we provide a comparison with existing pulsar observation sensitivities.
@@ I Introduction
Understanding dark matter remains one of the foremost questions in modern cosmology. An international effort combines theoretical, numerical, and observational efforts to search a dark matter parameter space covering over $\sim 100$ orders of decades of mass Feng (2010). At the lowest mass end of this spectrum, typically $m\lesssim 10^{-19}\,\mathrm{eV}$, wave-like phenomena manifest on astrophysical scales and impact large scale structure formation Hu et al. (2000). Moreover, ultraviolet-complete theories, such as string theory, predict a tower of light scalar particles arising from the compactification of extra dimensions Arvanitaki et al. (2010), making ultralight dark matter an appealing candidate.
@@
Originally, it was thought that a particle with $m\sim 10^{-22}\,\mathrm{eV}$ could potentially solve the core-cusp Navarro et al. (1996); Persic et al. (1996); Gentile et al. (2004), missing satellite Klypin et al. (1999); Moore et al. (1999), and too-big-to-fail problems Boylan-Kolchin et al. (2011) which described discrepancies between the predictions of dark matter only simulations and observations of small scale structure (see Weinberg et al. (2015); Bullock and Boylan-Kolchin (2017) for review) without invoking baryonic physics Hu et al. (2000). However, the originally proposed mass of $m\sim 10^{-22}\,\mathrm{eV}$ has since been ruled out by a wide range of phenomena including observations of: subhalo mass function ($>3\times 10^{-21}\,\mathrm{eV}$) Nadler et al. (2021); Schutz (2020), ultra-faint dwarf half-light radii ($>3\times 10^{-19}\,\mathrm{eV}$ Dalal and Kravtsov (2022); Marsh and Niemeyer (2019)) or cores Hayashi et al. (2021), galactic density profiles ($>10^{-20}\,\mathrm{eV}$ Bar et al. (2018, 2022)), satellite masses ($>6\times 10^{-22}\,\mathrm{eV}$ Safarzadeh and Spergel (2020)), Lyman-alpha forest ($>2\times 10^{-20}\,\mathrm{eV}$ Rogers and Peiris (2021)), strong lensing ($>4\times 10^{-21}\,\mathrm{eV}$ Powell et al. (2023)). For recent reviews, see Hui et al. (2017); Ferreira (2021); Hui (2021); Niemeyer (2020).
@@
Pulsar timing arrays (PTA) are thought to be sensitive to galactic metric perturbations Afzal et al. (2023); Antoniadis et al. (2022); Khmelnitsky and Rubakov (2014); Porayko et al. (2018); Chen et al. (2021); Goncharov et al. (2021); Verbiest et al. (2024)through precise measurements of pulses over $\mathcal{O}(10)$-year observations. This has already been studied in the context of gravitational waves backgrounds Agazie et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023), substructure of cold dark matter halo Baghram et al. (2011); Liu (2023), ultralight dark matter Compton fluctuations Khmelnitsky and Rubakov (2014); Porayko and Postnov (2014); Porayko et al. (2018); Smarra et al. (2023), and ultralight dark matter granule-star interactions Kim and Mitridate (2024). The effect of higher-spin ultralight dark matter fields on pulsar timing arrays has also been studied Armaleo et al. (2020); Liang and Trodden (2021); Zhang et al. (2023); Wu et al. (2023); Jain and Amin (2022a); Adshead and Lozanov (2021a). The gravitational redshift created by ultralight dark matter Compton fluctuations have been used to constrain the dark matter mass around $10^{-23}\,\mathrm{eV}$ Khmelnitsky and Rubakov (2014); Porayko et al. (2018); Smarra et al. (2023).
@@
In this paper, we study the Shapiro time delay and gravitational redshift caused by de Broglie scale oscillations in the density field. Interestingly, the granular structure oscillates on the time scales probed by experiments, $\sim\mathcal{O}(10)\,\mathrm{yrs}$, for field masses $\gtrsim 10^{-17}\,\mathrm{eV}$. This means that the oscillation of the granular structure may provide some of the highest mass probes of ultralight dark matter. Already de Broglie scale oscillations have been studied in the context of the dynamical heating of galactic stellar populations Dalal and Kravtsov (2022); Marsh and Niemeyer (2019); Chowdhury et al. (2023), relativistic Doppler shifts in pulsars Kim and Mitridate (2024), and astrometry Kim (2024); Dror and Verner (2024).
@@
We provide simulations of mock pulsars in an oscillating ultralight dark matter background. We track the Shapiro time delay and gravitational redshift between an observer and the pulsars over time and compare simulation results to analytic predictions. We characterize these time delay signals by calculating the root-mean-square amplitude, temporal power spectrum, and dependence on dark matter and pulsar-observer parameters. Finally, we compare the result of this work to the existing ultralight dark-matter constraints from pulsar timing arrays, and estimate the sensitivity of current PTA experiments to these granular phenomena. We conclude that both the Shapiro time delay and gravitational redshift effect of the granules would produce unique temporal power spectra in pulsar timing array systems, which, if observed, would provide a smoking gun for ultralight dark matter. However, current experiments are unlikely to be sensitive to this effect.
@@ II.1 Ultra light dark matter
Ultralight dark matter is usually treated as a classical Eberhardt et al. (2024) and non-relativistic wavelike field. In this paper, we will assume that it is a scalar field, but generalization of higher spins also exists Jain and Amin (2022b); Adshead and Lozanov (2021b). We start with the Klein-Gordon equation for a quadratic potential
@@
and relates the Newtonian potential, $\Phi$ to the field amplitude. Here we have chosen the field norm such that $|\psi|^{2}=\rho$ is the spatial density field.
@@
$\lambda_{\mathrm{db}}$ $=2\pi\hbar/m\sigma\sim 6\times 10^{-6}\left(\frac{10^{-17}\,\mathrm{eV}}{m}\right)\,\mathrm{kpc}\,.$ (7)
Where in this work $\sigma$ is the velocity dispersion of the dark matter halo. In the following work, it will generally be the velocity dispersion of the system.
@@
$\psi(\vec{r})$ $=\braket{\rho(}{\vec{R}+\vec{r}}{)}\int d^{3}\vec{v}\sqrt{f(v)}\,e^{-im\vec{v}\cdot\vec{r}/\hbar}\,,$ (8)
@@
where $\sigma$ is the dark matter velocity dispersion at $\vec{R}$. This holds as long as we only look over scales that are smaller than the galactic halo scale radius, i.e. $r\lesssim R_{s}$. We will take the typical Earth-pulsar separation to be $\sim 1\,\mathrm{kpc}$ which is smaller than the scale radius of the Milky-way halo ($R_{s}\sim 2\,\mathrm{kpc}$ Porcel et al. (1997)).
The interference patterns create granule structures in the density. It is often a useful approximation to treat these granules as quasi-particles with an effective radius approximately equal to the de Broglie wavelength $r_{\mathrm{eff}}\sim\lambda_{\mathrm{db}}$, and effective mass $m_{\mathrm{eff}}\sim\lambda_{\mathrm{db}}^{3}\rho$ Dalal and Kravtsov (2022); Kim and Mitridate (2024). The granules persist for approximately the crossing time of the granule which is given by the de Broglie time $\tau_{\mathrm{db}}$.
@@ II.2 Gravitational time delays
$ds^{2}=\left(1+2\frac{\Phi}{c^{2}}\right)\,dt^{2}-\left(1-2\frac{\Psi}{c^{2}}\right)\,dx^{2}\,,$ (10)
where we will also assume that $\Phi=\Psi$ is just the Newtonian potential calculated in (4). This potential then creates a redshift, at time $t$, in a pulsar with frequency $\Omega_{p}$ at position $r_{p}$, as observed by an observer at $r_{e}$ given
@@
where $\hat{n}_{i}=(\vec{r}_{e}-\vec{r}_{p})/|(\vec{r}_{e}-\vec{r}_{p})|$ is the unit vector in the direction of signal propagation and $x(t^{\prime\prime})$ is the path of the signal from the pulsar to observer. $t$ is the time the pulse is observed at Earth and $t^{\prime}$ is the time it is emitted from the pulsar. The first line in equation (11) gives the gravitational redshift of the frequency and the second the Shapiro time delay shift in the frequency. The effects then given a time delay via integration
[…]
@@ III.1 Initial conditions
with $\sigma$ the velocity dispersion of the dark matter. We will take $\sigma=200\,\mathrm{km/s}$. The resulting density pattern can be seen in the left panel of Figure 2. The power spectrum of fluctuations created by these initial conditions consistently models the fluctuations around the average density profile in typical ultralight dark matter halos Hui et al. (2021); Dalal et al. (2021), see Appendix 10.
@@ III.2 Solver
where $\mathcal{F}$ is a discrete fast Fourier transform.
III.3 Pulsars
Figure 3: The Shapiro time delay (left) and gravitational redshift (right) for a simulated pulsar at $1\,\mathrm{kpc}$ from Earth in a ULDM density field at the local dark matter density. Note that the mean is removed from each set of data as it would be indistinguishable from the intrinsic pulsar frequency. The mass of the dark matter field in this simulation was $m=10^{-22}\,\mathrm{eV}$.
Pulsar and observer locations are chosen within the simulation volume, $r_{p}$ and $r_{e}$ respectively. The Shapiro time delay and gravitational redshift between each pulsar and Earth are calculated, at each timestep, $t_{i}$, see Figure 3. The Shapiro time delay between Earth and the $j$th pulsar at time $t_{i}$ is given
$\delta t_{j}^{sh}(t_{i})=\int_{r^{j}_{p}}^{r_{e}}\frac{dl}{c}\,\frac{\Phi(l,t_{i})}{c^{2}}\,,$ (16)
where $r^{j}_{p}$ is the location of the $j$th pulsar. The integrand is plotted for an example simulation realization in the right panel of Figure 2.
@@ IV Results
Figure 4: The average Shapiro time delay measured for pulsars at a distance of $1\,\mathrm{kpc}$ (left) and the average gravitational potential fluctuations (right), $\braket{\Phi}_{\mathrm{rms}}$, for a set of plane-wave box simulations, described in section III.1, as a function of the field mass $m_{22}$. The data points for the Shapiro delays correspond to the average over all the pulsars at a single moment in time. The large scatter around the prediction is due to the random potential fluctuations around the observer at that time in each simulation. The quasi-particle approximation, equation (23), is shown in red. The quasi-particle approximation provides a reasonably accurate estimation of time delays and potential fluctuations.
@@
Throughout this section, we will approximate quantities using the quasi-pariticle approximation and then compare these approximations to the results of simulated data. In this approximation we assume that the density field can be described by a collection of tightly packed “quasi-particles” modeled as spheres with radius equal to the de Broglie wavelength, $\lambda_{\mathrm{db}}$, and effective mass $m_{\mathrm{eff}}\sim\rho\lambda^{3}_{\mathrm{db}}$, with $\rho$ the local field density. In the context of ultralight dark matter, this approximation has been previously applied to the description of the dynamical heating of dwarf galaxies Dalal and Kravtsov (2022); Zupancic and Widrow (2023), pulsar Doppler shifts Kim and Mitridate (2024), and astrometry Kim (2024).

### v2

@@ Abstract
The study of ultralight dark matter helps to constrain the lower bound of the mass in minimally coupled dark matter models. The granular structure of ultralight dark matter density fields produces metric perturbations which have been identified as a potentially interesting probe of this model. For dark matter masses $m\gtrsim 10^{-17}\,\mathrm{eV}$, these perturbations would fluctuate on timescales comparable to observational timescales. In this paper, we estimate the expected time delay these fluctuations would generate in simulated pulsar signals. We simulate arrays of mock pulsars in a fluctuating granular density field. We calculate the expected Shapiro time delay, gravitational redshift, and Doppler shift and compare analytical estimates with the results of simulations. Finally, we provide a comparison with existing pulsar observation sensitivities.
@@ I Introduction
Understanding dark matter remains one of the foremost questions in modern cosmology. An international effort combines theoretical, numerical, and observational efforts to search a dark matter parameter space covering over $\sim 80$ orders of decades of mass Fen (2010). At the lowest mass end of this spectrum, typically $m\lesssim 10^{-19}\,\mathrm{eV}$, wave-like phenomena manifest on astrophysical scales and impact large scale structure formation Hu et al. (2000). Moreover, ultraviolet-complete theories, such as string theory, predict a tower of light scalar particles arising from the compactification of extra dimensions Arvanitaki et al. (2010), making ultralight dark matter an appealing candidate.
@@
Originally, it was thought that a particle with $m\sim 10^{-22}\,\mathrm{eV}$ could potentially solve the core-cusp Navarro et al. (1996); Persic et al. (1996); Gentile et al. (2004), missing satellite Klypin et al. (1999); Moore et al. (1999), and too-big-to-fail problems Boylan-Kolchin et al. (2011) which described discrepancies between the predictions of dark matter only simulations and observations of small scale structure (see Weinberg et al. (2015); Bullock and Boylan-Kolchin (2017) for review) without invoking baryonic physics Hu et al. (2000). However, the originally proposed mass of $m\sim 10^{-22}\,\mathrm{eV}$ has since been ruled out by a wide range of phenomena including observations of: subhalo mass function ($3\times 10^{-21}\,\mathrm{eV}$) Nadler et al. (2021); Schutz (2020), ultra-faint dwarf half-light radii ($3\times 10^{-19}\,\mathrm{eV}$ Dalal and Kravtsov (2022); Marsh and Niemeyer (2019)) or cores Hayashi et al. (2021), galactic density profiles ($10^{-20}\,\mathrm{eV}$ Bar et al. (2018, 2022)), satellite masses ($6\times 10^{-22}\,\mathrm{eV}$ Safarzadeh and Spergel (2020)), Lyman-alpha forest ($2\times 10^{-20}\,\mathrm{eV}$ Rogers and Peiris (2021)), strong lensing ($4\times 10^{-21}\,\mathrm{eV}$ Powell et al. (2023)). For recent reviews, see Ferreira (2021); Hui (2021); Niemeyer (2020); Eberhardt and Ferreira (2025).
@@
Pulsar timing arrays (PTA) are thought to be sensitive to galactic metric perturbations Afzal et al. (2023); Antoniadis et al. (2022); Khmelnitsky and Rubakov (2014); Porayko et al. (2018); Chen et al. (2021); Goncharov et al. (2021); Verbiest et al. (2024)through precise measurements of pulses over $\mathcal{O}(10)$-year observations. This has already been studied in the context of gravitational waves backgrounds Agazie et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023), substructure of cold dark matter halo Baghram et al. (2011); Liu (2023), primordial black holes Dror et al. (2019), ultralight dark matter Compton fluctuations Khmelnitsky and Rubakov (2014); Porayko and Postnov (2014); Porayko et al. (2018); Smarra et al. (2023), and ultralight dark matter granule-star interactions Kim and Mitridate (2024). The effect of higher-spin ultralight dark matter fields on pulsar timing arrays has also been studied Armaleo et al. (2020); Liang and Trodden (2021); Zhang et al. (2023); Wu et al. (2023); Jain and Amin (2022); Adshead and Lozanov (2021a); Chowdhury et al. (2024). The gravitational redshift created by ultralight dark matter Compton fluctuations have been used to constrain the dark matter mass around $10^{-23}\,\mathrm{eV}$ Khmelnitsky and Rubakov (2014); Porayko et al. (2018); Smarra et al. (2023). Likewise, the impact of ULDM on the frequency of gravitational waves has been studied in Brax et al. (2024); Blas et al. (2025).
@@
In this paper, we study the Shapiro time delay, gravitational redshift, and Doppler shift caused by de Broglie scale oscillations in the density field. Interestingly, the granular structure oscillates on the time scales probed by experiments, $\sim\mathcal{O}(10)\,\mathrm{yrs}$, for field masses $\gtrsim 10^{-17}\,\mathrm{eV}$. This means that the oscillation of the granular structure may provide some of the highest mass probes of ultralight dark matter. Already de Broglie scale oscillations have been studied in the context of the dynamical heating of galactic stellar populations Dalal and Kravtsov (2022); Marsh and Niemeyer (2019); Chowdhury et al. (2023), relativistic Doppler shifts in pulsars Kim and Mitridate (2024), stochastic lensing Eberhardt et al. (2025), and astrometry Kim (2024); Dror and Verner (2024).
@@
We provide simulations of mock pulsars in an oscillating ultralight dark matter background. We track the Shapiro time delay, gravitational redshift, and Doppler shift between an observer and the pulsars over time and compare simulation results to analytic predictions. We characterize these time delay signals by calculating the root-mean-square amplitude, and dependence on dark matter and pulsar-observer parameters. Finally, we compare the result of this work to the existing ultralight dark-matter constraints from pulsar timing arrays, and estimate the sensitivity of current PTA experiments to these granular phenomena. We conclude that both the Shapiro time delay and gravitational redshift effect of the granules would produce unique temporal power spectra in pulsar timing array systems, which, if observed, would provide a smoking gun for ultralight dark matter. Our simulations of the Doppler shift corroborate the conclusions of Kim and Mitridate (2024). However, current experiments are unlikely to be sensitive to any of these effects.
@@ II.1 Ultra light dark matter
Ultralight dark matter is usually treated as a classical Eberhardt et al. (2024) and non-relativistic wavelike field. In this paper, we will assume that it is a scalar field, but generalization of higher spins also exists Jain and Amin (2022); Adshead and Lozanov (2021b). We start with the Klein-Gordon equation for a quadratic potential
@@
and relates the Newtonian potential, $\Phi$ to the field amplitude. Here we have chosen the field norm so that $|\psi|^{2}=\rho$ is the spatial density field.
@@
$\lambda_{\mathrm{db}}$ $=2\pi\hbar/m\sigma\sim 6\times 10^{-6}\left(\frac{10^{-17}\,\mathrm{eV}}{m}\right)\left(\frac{200\,\mathrm{km/s}}{\sigma}\right)\,\mathrm{kpc}\,.$ (7)
Where in this work $\sigma$ is the velocity dispersion of the dark matter halo. In the following work, $\sigma$ will generally refer to the the velocity dispersion of the system. We note here that the definition of these time and length scales often differs in the literature up to a factor of $2\pi$, for example our notation differs from Kim and Mitridate (2024). However, keeping definitions consistent, we will recover similar results.
@@
$\psi(\vec{r})$ $=\sqrt{\braket{\rho(|\vec{R}+\vec{r}|)}}\int d^{3}\vec{v}\sqrt{f(v)}\,e^{-im\vec{v}\cdot\vec{r}/\hbar}\,$, (8)
@@
where $\sigma$ is the dark matter velocity dispersion at $\vec{R}$. This holds as long as we only look over scales that are smaller than the galactic halo scale radius, i.e. $r\lesssim R_{s}$. We will take the typical Earth-pulsar separation to be $\sim 1\,\mathrm{kpc}$, which is smaller than the scale radius of the Milky-way halo ($R_{s}\sim 2\,\mathrm{kpc}$ Porcel et al. (1997)).
The interference patterns create granule structures in the density. It is often a useful approximation to treat these granules as quasiparticles with an effective radius approximately equal to the de Broglie wavelength $r_{\mathrm{eff}}\sim\lambda_{\mathrm{db}}$, and effective mass $m_{\mathrm{eff}}\sim\lambda_{\mathrm{db}}^{3}\rho$ Dalal and Kravtsov (2022); Kim and Mitridate (2024). The granules persist for approximately the crossing time of the granule which is given by the de Broglie time $\tau_{\mathrm{db}}$.
@@ II.2 Gravitational time delays
$ds^{2}=\left(1+2\frac{\Phi}{c^{2}}\right)\,c^{2}dt^{2}-\left(1-2\frac{\Psi}{c^{2}}\right)\,dx^{2}\,$, (10)
where we will also assume that $\Phi=\Psi$ is just the Newtonian potential calculated in (4).
II.2.1 Redshift and Shapiro delay
This potential then creates a redshift, at time $t$, in a pulsar with frequency $\Omega_{p}$ at position $r_{p}$, as observed by an observer at $r_{e}$ given
@@ II.2.1 Redshift and Shapiro delay
where $\hat{n}_{i}=(\vec{r}_{e}-\vec{r}_{p})/|(\vec{r}_{e}-\vec{r}_{p})|$ is the unit vector in the direction of signal propagation and $x(t^{\prime\prime})$ is the path of the signal from the pulsar to observer. $t$ is the time the pulse is observed at Earth, and $t^{\prime}$ is the time it is emitted from the pulsar. The first line in equation (11) gives the gravitational redshift of the frequency, and the second the Shapiro time delay shift in the frequency. The effects then given a time delay via integration
$\delta t(T)=\int_{0}^{T}\frac{\Delta\Omega_{p}(t^{\prime})}{\Omega_{p}}dt^{\prime}\,$, (12)
@@
II.2.2 Doppler shift
The fluctuating potential gradients also impact the proper motion between observer and pulsar. This relative proper motion results in a Doppler shift of the observed frequency of the pulsar. This shift can be found by integrating the relative forces due to the gravitational potential, $\Phi$, giving a frequency shift due to the Doppler effect of
$z_{D}=\int^{t}\hat{n}_{i}\partial_{i}\left(\frac{\Phi(t^{\prime}-D/c,\,r_{p})-\Phi(t^{\prime},\,r_{e})}{c}\right)\,dt^{\prime}.$ (13)
Where $D$ is the distance between the pulsar and the observer, approximated as constant over the observing time $t$.
@@ III.1 Initial conditions
$\sigma$ is the velocity dispersion of the dark matter. We will take $\sigma=200\,\mathrm{km/s}$. The resulting density pattern can be seen in the left panel of Figure 2. The power spectrum of fluctuations created by these initial conditions consistently models the fluctuations around the average density profile in typical ultralight dark matter halos Hui et al. (2021); Dalal et al. (2021), see Appendix 11.
@@ III.2 Solver
where $\mathcal{F}$ is a discrete fast Fourier transform. We note that because the system does not evolve under its own gravitational potential that the time step $\Delta t$ can be chosen arbitrarily. In principle, we could integrate equation (14) directly and then reconstruct the field at any time by re-summing these plane-waves. This method would have a similar runtime depending on the simulation parameters chosen.
III.3 Gravitational redshift and Shapiro delay
Figure 3: The Shapiro time delay (left), gravitational redshift (center), and Doppler shift (right) for a simulated pulsar at $1\,\mathrm{kpc}$ from Earth in a ULDM density field at the local dark matter density. The mass of the dark matter field in this simulation was $m=10^{-22}\,\mathrm{eV}$.
Pulsar and observer locations are chosen within the simulation volume, $r_{p}$ and $r_{e}$ respectively. The Shapiro time delay and gravitational redshift between each pulsar and Earth are calculated, at each timestep, $t_{i}$, see Figure 3. The expected root-mean-square Shapiro time delay between Earth and the $j$th pulsar at time $t_{i}$ can be estimated
$\braket{\delta t_{j}^{sh}(t_{i})}_{\mathrm{rms}}=\int_{r^{j}_{p}}^{r_{e}}\frac{dl}{c}\,\frac{\delta\Phi(l,t_{i})}{c^{2}}\,$, (17)
where $r^{j}_{p}$ is the location of the $j$th pulsar. This approximation is discussed in more detail in Appendix A. The integrand is plotted for an example simulation realization in the right panel of Figure 2.
@@ III.3 Gravitational redshift and Shapiro delay
III.4 Doppler shift
To model the Doppler shift we need to track the relative motion of pulsars not just the gravitational potential. Therefore, in these simulations we treat the pulsars as an ensemble of massless point particles with uniform random initial positions in the box and velocities drawn from the same Maxwell-Boltzmann distribution as the dark matter, i.e., drawn from distribution in equation (15). The point particles are coupled to the gravitational potential of the ultralight dark matter and updated using a drift-kick-drift scheme where the drift half-step now involves both an update of the ultralight dark matter field and the $i$th particle as
$\psi(t+\Delta t/2)$ $=\mathcal{F}^{-1}\left[e^{-i\hbar k^{2}\Delta t/4m}\mathcal{F}\left[\psi_{i}(t)\right]\right]\,$, (21)
$\vec{r}_{i}(t+\Delta t/2)$ $=\vec{r}_{i}+\vec{v}_{i}(t)\,\Delta t/2\,$, (22)
and a kick full-step as
$\vec{v}_{i}(t+\Delta t)=\vec{v}_{i}+\nabla\phi(\vec{r}_{i})\,\Delta t\,.$ (23)
Where the potential gradient is calculated using a finite (5-pt) differentiation stencil and linearly interpolated at the particles position.
The Doppler shift for the $i$th particle at each data drop can be calculated using the velocity along the line of sight to the observer as
$z_{D}^{i}(t)=\frac{\hat{n}_{i}\cdot\vec{v}_{i}}{c}\,.$ (24)
@@ IV Results
Figure 4: The average Shapiro time delay measured for pulsars at a distance of $1\,\mathrm{kpc}$ (left) and the average gravitational potential fluctuations (center), $\braket{\Phi}_{\mathrm{rms}}$, the average Doppler shift per root time (right), for a set of plane-wave box simulations, described in section III.1, as a function of the field mass $m_{22}$. The data points for the Shapiro delays correspond to the average over all the pulsars at a single moment in time. The large scatter around the prediction is due to the random potential fluctuations around the observer at that time in each simulation. The quasi-particle approximation for each effect equation is shown in red. For Shapiro delays the prediction is equation (26), for the potential rms the prediction is given by equation (28), and for the Doppler shifts the prediction is given by equation (33). The quasi-particle approximation provides a reasonably accurate estimation in each situation.
@@
Throughout this section, we will approximate quantities using the quasi-pariticle approximation and then compare these approximations to the results of simulated data. In this approximation we assume that the density field can be described by a collection of tightly packed “quasi-particles” modeled as spheres with radius equal to the de Broglie wavelength, $\lambda_{\mathrm{db}}$, and effective mass $m_{\mathrm{eff}}\sim\rho\lambda^{3}_{\mathrm{db}}$, with $\rho$ the local field density. In the context of ultralight dark matter, this approximation has been previously applied to the description of the dynamical heating of dwarf galaxies Dalal and Kravtsov (2022), pulsar Doppler shifts Kim and Mitridate (2024), and astrometry Kim (2024).